Job Costs, Ending Work in Process During March, Aragon Company worked on three jobs.

Data relating to these three jobs follow:

Job 78 Job 79 Job 80

Units in each order 120 210 230

Units sold — 210 —

Materials requisitioned $912 $945 $2,553

Direct labor hours 132 231 391

Direct labor cost $1,584 $3,927 $3,206

Overhead is assigned on the basis of direct labor hours at a rate of $7.50 per direct labor hour. During March, Jobs 78 and 79 were completed and transferred to Finished Goods Inventory. Job 79 was sold by the end of the month. Job 80 was the only unfinished job at the end of the month.

Required: 1. Calculate the per-unit cost of Jobs 78 and 79. Round your answers to the nearest cent. Job 78 $ per unit Job 79 $ per unit

2. Compute the ending balance in the work-in-process inventory account. Round your answer to the nearest cent.

3. Prepare the journal entries reflecting the completion of Jobs 78 and 79 and the sale of Job 79. The selling price is 130 percent of cost. Round your answers to the nearest cent.
Answer #1

1)

Job 78 Job 79
Direct material 912 945
Direct labor 1584 3927
Overhead applied 990     [7.5*132] 1732.5     [231*7.5]
Total cost 3486 6604.5
Unit cost [Total cost /Number of units] 3486 /120= 29.05 6604.5/210= 31.45

2)

Cost of ending work in process is equal to cost of job 80 : 2553 Direct material +3206 direct labor +(391*7.5) overhead

                          = 2553 +3206 +2932.5

                      = 8691.50

.3)

Date Account title Debit credit
1 Finished goods inventory 10090.5
Work in process 10090.5
[Being Job 78 and 79 completed and transferred out 3486+6604.5]
2 Cash /Accounts receivable (6604.5 *130%) 8585.85
sales revenue 8585.85
[being sales made]
cost of goods sold 6604.5
Finished goods inventory 6604.5